Transportation Method (Example) Mar 800 Demand Capacity: 1000 1,100 Regular Overtime Subcontracting 800 100 200 100 Costs 700 100 300 700 50 150 Beginning invento $30 per tire $50 per tire Regular time Overtime Subcontracting$60 per tire Carrying $4 per tire per month

We use the transportation tableau to find the optimal solution

please show working
Answer #1

In order to formulate the problem as a transportation model, we need to create the parameter table. Rows of the parameter table show supply period and columns indicate demand period. Regular cost of production in a supply period to fulfil the demand of same period is as given in the problem. For example regular production cost in period 1 to fulfil the demand of period 1 is $ 30 and overtime cost in the same supply and demand period is $ 50. Subcontract cost is $ 60.

Regular time Cost of fulfilling demand of n months later than the current month increases by n times the carrying cost. For example, cost of fulfilling demand of month 3 (2 periods later) from overtime time supply of month 1 = 50+3*4 = 62

While calculating net demand, opening inventory is subtracted from first month demand and required closing inventory is added to the last month demand.

Q fx -SUMPRODUCT(B11:D19,H11:J19) SSolver Parameters 1Demand & Supply Mth1 Mth 2 Mth 3 2 Demand 3 Capac 4 Regular 5 Overtime 6 Subcontractin 7 Beginning Invento Cost data Regular time cost $30 Overtime cost Subcontracting cost $60 1,000 1,100800 Set Objective: SKS8 $50 O Max Min Value Of:o 800 100 700 100 300 700 By Changing Variable Cells: ing Cost HS11:SJS13,SIS14:SJS16,SJS17:SJS19 Subject to the Constraints: SH$20:SIS20 SB$20:SDS20 SK$11:SKS19SES11 SES19 150 100 Add Total Cost 99,500 9 Parameter Table Optimal production schedule Change Mth 1 Mth 2 Mth 3 Su Mth 1 Mth 2 Mth 3 Row Total 11 Mth 1 Regular 12 Mth 1 Overtime 13 Mth 1 Subcontract 14 Mth 2 Regular 15 Mth 2 Overtime 16 Mth 2 Subcontradt 17 Mth 3 Regular 18 Mth 3 Overtime 19 Mth 3 Subcontract 20 Net Demand Mth 1 Regular Mth 1 Overtime Mth 1 Subcontract Mth 2 Regular Mth 2 Overtime Mth 2 Subcontract Mth 3 Regular Mth 3 Overtime Mth 3 Subcontract Column Total 800 100 200 Delete 54 100 Reset All 34 Make Unconstrained Variables Non-Negative 100 300 100 100 300 54 Select a Solving Method: Simplex LP Options 30 700 50 50 Solving Method Select the IPOPT Nonlinear engine for Solver Problems that are smooth nonlinear.Select the LP Simplex engine for linear Solver Problems 150 50 900 1,100800 900 1,100 800 Solvelose

EXCEL Formulas:

Cell Formula Copy to
K11 =SUM(H11:J11) K11:K19
H20 =SUM(H11:H19) H20:J20
K8 =SUMPRODUCT(B11:D19,H11:J19)

Total cost = $ 99,500